Eastern Motors Auto Dealership wanted to estimate the average CLV over a 5 year time horizon of a customer who purchases a new vehicle. The average vehicle sells for $25,500 and has a margin of 12%. Based on historical averages, 74% of people buying a new vehicle at Eastern will return for service 11 times over the next 5 years. Though it varies considerably, Eastern generates approximately $118 in margin on each service visit after accounting for parts and direct labor costs. What is the total estimated CLV over a 5 year time horizon for someone who purchases a new vehicle at Eastern Motors?
Answer #1

Margin on selling only vehicle= 25500*12%=$3060

For 74% of customer who comes for services generates margin over 5 years= 11*118=$1298

So, the average CLV= 3060+0.74*1298+0.26*0=$4020.52